Industry A has four firms. The largest firm in Industry A has more than 90 percent of the market share. Industry B also has four firms, but each of those four firms in Industry B has 25 percent of the market share. The Herfindahl-Hirschman index will be 


A. larger for Industry B than Industry A, but the four-firm concentration will be the same. 

В. tte same for both industries, but the four-firm concentration will be larger for industry Y than Industry A 

C. larger for Industry Athan Industry B, but the four-firm concentration will be the same. 

D. the same for both industries, but the four-firm concentration will be larger for industry Athan Industry B.
Answer #1

Industry A has 4 firms, one firm has 90% market share. Let us assume other firms Market share is 4%,3%,3%.

HHI of industry A = 902 + 42 + 32 + 32 = 8100+16+9+9 = 8134

Industry B has four firms with equal share.

HHI = 252 + 252 + 252 + 252 = 625+625+625+625= 2500

Thus industry A has high HHI.

C. Larger for industry A but the four firm concentration would be same.

The firm concentration would be 100% in A as well as B. Since, each industry has only 4 firms.

Please contact if having any query thank you.